Adam likes his caffé latte prepared to contain exactly 1/4 espresso and 3/4 steamed milk by volume. On a graph with the volume of steamed milk on the horizontal axis and the volume of espresso on the vertical axis, draw two of his indifference curves, U1 and U2, with U1 > U2.
Answer #1

Since each part of caffé latte has 0.25 part of espresso ES and 0.75 part of steamed milk SM, the preferences indicate a perfect complementary relationship between espresso and steamed milk. This implies, ES/SM = 0.25/0.75 or 3ES = SM.

The utility function is given by U = min (3ES, SM). We have U1 = 9 so that SM = 9 and ES = 3. Similarly, U2 = 2 where SM = 2 and ES = 0.67

Volume of espressd 0.67 Volume of steamed milk